### v1

# On polynomial automorphisms of Nagata type

Abstract.
We define a family of polynomial ring homomorphisms generalising the Nagata automorphism, and relate the condition of non-vanishing of their Jacobian determinant with a certain linear partial differential equation. We also give a condition on that family to get wild automorphisms.
Key words and phrases:
polynomial automorphisms, Nagata automorphism, wild automorphisms, linear PDE
2010 Mathematics Subject Classification:
14R10, 13B25, 13F20, 35F05
1. Introduction
In what follows $F$ will be a field of characteristic 00 unless otherwise stated. The polynomials:
$f(x,y,z)$ $=x-2y(zx+y^{2})-z(zx+y^{2})^{2},$
$g(x,y,z)$ $=y+z(zx+y^{2}),$
$h(x,y,z)$ $=z.$
define an automorphism on $F[x,y,z]$ called Nagata automorphism [1]. The determinant of the Jacobian matrix of this automorphism es $1$. In the following paragraph we recall the notion of tame automorphisms of polynomials. Let $1\leq i\leq n$, the $i$th elementary automorphism $\Phi^{(i)}\colon F^{n}\to F^{n}$ is the automorphism defined as
$\Phi^{(i)}(x_{1},\dots,x_{n})=\Big{(}x_{1},\dots,x_{i-1},x_{i}+f(x_{1},\dots,x_{i-1},x_{i+1},\dots,x_{n}),x_{i+1},\dots,x_{n}\Big{)}$
and its inverse is the automorphism
${\Phi^{(i)}}^{-1}(x_{1},\dots,x_{n})=\Big{(}x_{1},\dots,x_{i-1},x_{i}-f(x_{1},\dots,x_{i-1},x_{i+1},\dots,x_{n}),x_{i+1},\dots,x_{n}\Big{)}\,.$
One may notice that the Jacobian matriz of $\Phi^{(i)}$ has determinant $1$. An automorphism is tame if it is a finite composition of elementary and linear automorphisms. An automorphism that is not tame is called wild. In 1972, Nagata conjectured that the Nagata automorphism is wild. In [2, Corollary 9], the authors proved this conjecture. In this work, we define ring homomorphisms on $F[x,y,z]$, given in (2), extending the Nagata automorphism which are wild automorphisms under some assumptions, see Corollary 1). We also relate the condition of non-vanishing of their Jacobian determinant with a certain linear partial differential equation, see Theorem 1. As a consequence, we prove that those ring homomorphisms on $F[x,y,z]$ with nonzero Jacobian determinant, are in fact automorphisms, see Theorem 2.
2. Nagata polynomial map
For a polynomial $\varphi\in F[x,y,z]$, we set
(1) $\begin{split}f(x,y,z)&=x-2y\varphi-z\varphi^{2}\,,\\ g(x,y,z)&=y+z\varphi\,,\\ h(x,y,z)&=z\,.\end{split}$
The triplet $(f,g,h)$ induces a ring homomorphism on $F[x,y,z]$. Note that $(f,g,h)$ is the identity when one takes $\varphi=0$, so $(f,g,h)$ may be seen as a perturbation of the identity. If $\varphi=xz+y^{2}$ then $(f,g,h)$ is the Nagata automorphism. The Milnor number of $(f,g,h)$ is the dimension
$\mu(f,g,h)=\dim_{F}\frac{F[x,y,z]}{\langle f,g,h\rangle}$
as vector space over $F$.
Lemma 1.
If $(f,g,h)$ is a ring homomorphism as in (2), then $\mu(f,g,h)=1$.
Proof.
One has,
$y$ $=(y+z\varphi)-\varphi z=g-\varphi z,$
$x$ $=(x-2y\varphi-z\varphi^{2})+2(y+z\varphi)\varphi-\varphi^{2}z=f+2\varphi g-\varphi^{2}z,$
hence we get an equality $\langle f,g,h\rangle=\langle x,y,z\rangle$ of polynomial ideals, and this implies that $\mu(f,g,h)=1$. ∎
Remark 1.
Note that $(f,g,h)$ is not always an automorphism for any $\varphi$ despite the fact that $\mu(f,g,h)=1$ by Lemma 1. For instance, if $\varphi=x$ or $\varphi=y$, then $(f,g,h)$ is not an automorphism. In the first case, $(f,g,h)$ sends both points $(0,0,1)$ and $(-1,1,1)$ to $(0,0,1)$. Theorem 2 provides a necessary and sufficient condition on $\varphi$ such that $(f,g,h)$ is an automorphism.
Proposition 1.
The Jacobian matrix $J(f,g,h)$ has nonzero determinant in $F$ if and only if $\varphi$ satisfies the linear PDE:
$-2y\varphi_{x}+z\varphi_{y}=0\,.$
Proof.
We have
$f_{x}$ $=1-2y\varphi_{x}-2z\varphi\varphi_{x},$
$f_{y}$ $=-2\varphi_{x}-2y\varphi_{y}-2z\varphi\varphi_{y},$
$g_{x}$ $=z\varphi_{x},$
$g_{y}$ $=1+z\varphi_{y}\,.$
Hence the Jacobian matrix $J(f,g,h)$ is
$\begin{pmatrix}1-2y\varphi_{x}-2z\varphi\varphi_{x}&-2\varphi_{x}-2y\varphi_{y}-2z\varphi\varphi_{y}&f_{z}\\ z\varphi_{x}&1+z\varphi_{y}&g_{z}\\ 0&0&1\end{pmatrix}$
whose determinant is
$\begin{vmatrix}1-2y\varphi_{x}-2z\varphi\varphi_{x}&-2\varphi_{x}-2y\varphi_{y}-2z\varphi\varphi_{y}\\ z\varphi_{x}&1+z\varphi_{y}\end{vmatrix}=1-2y\varphi_{x}+z\varphi_{y}\,.$
If $-2y\varphi_{x}+z\varphi_{y}=0$ then the determinant of $J(f,g,h)$ is $1$. Conversely, if the determinant of $J(f,g,h)$ is a constant $c\in F-\{0\}$, then
$1-2y\varphi_{x}+z\varphi_{y}=c\,.$
However this equality is only possible when $c=1$, therefore $-2y\varphi_{x}+z\varphi_{y}=0$. ∎
Proposition 2.
Let $p\in F[t_{1},t_{2}]$ and set $\varphi=p(xz+y^{2},z)$ in $F[x,y,z]$. Then the ring homomorphism $(f,g,h)$ given in (2) is an automorphism on $F[x,y,z]$ with inverse:
(2) $\begin{split}f^{\prime}(x,y,z)&=x+2y\varphi-z\varphi^{2},\\ g^{\prime}(x,y,z)&=y-z\varphi,\\ h^{\prime}(x,y,z)&=z\,.\end{split}$
Proof.
We have
$\varphi(f^{\prime},g^{\prime},h^{\prime})$ $=p(f^{\prime}z+g^{\prime 2},z)$
$=p\Big{(}(x+2y\varphi-z\varphi^{2})z+(y-z\varphi)^{2},z\Big{)}$
$=p(xz+y^{2},z)$
$=\varphi\,,$
thus $\varphi(f^{\prime},g^{\prime},h^{\prime})=\varphi$. Hence,
$f(f^{\prime},g^{\prime},h^{\prime})$ $=f^{\prime}-2g^{\prime}\varphi(f^{\prime},g^{\prime},h^{\prime})-z\varphi^{2}(f^{\prime},g^{\prime},h^{\prime})$
$=(x+2y\varphi-z\varphi^{2})-2(y-z\varphi)\varphi-z\varphi^{2}$
$=x\,.$
On the other hand, we have
$g(f^{\prime},g^{\prime},h^{\prime})$ $=g^{\prime}+z\varphi(f^{\prime},g^{\prime},h^{\prime})$
$=(y-z\varphi)+z\varphi$
$=y\,.$
Therefore, the composite $(f,g,h)\circ(f^{\prime},g^{\prime},h^{\prime})$ is the identity. Similarly, one verifies that $(f^{\prime},g^{\prime},h^{\prime})\circ(f,g,h)$ is also the identity. ∎
For a polynomial $\psi$, we denote by $\overline{\psi}$ the highest degree homogeneous component of $\psi$.
Proposition 3.
Let $\varphi\in F[x,y,z]$ and suppose that $(f,g,h)$ is an automorphism on $F[x,y,z]$ of the form (2). If one the partial derivatives $\overline{\varphi}_{x},\overline{\varphi}_{y}$ is nonzero, then the automorphism $(f,g,h)$ is wild.
Proof.
Let $\overline{f},\overline{g},\overline{h}$ be the highest homogeneous component of $f,g,h$, respectively. Since $\overline{\varphi}_{x}$ or $\overline{\varphi}_{y}$ is nonzero, we have $\deg(\overline{\varphi})>0$, so
$\begin{split}\overline{f}&=-z\overline{\varphi}^{2}\,,\\ \overline{g}&=z\overline{\varphi}\,,\\ \overline{h}&=z\,.\end{split}$
Hence,
$\overline{f}_{x}\overline{g}_{y}-\overline{f}_{y}\overline{g}_{x}$ $=0$
$\overline{f}_{y}\overline{g}_{z}-\overline{f}_{z}\overline{g}_{y}$ $=-z\,\overline{\varphi}^{2}\overline{\varphi}_{y}$
$\overline{f}_{x}\overline{g}_{z}-\overline{f}_{z}\overline{g}_{x}$ $=-z\,\overline{\varphi}^{2}\overline{\varphi}_{x}$
$\overline{g}_{x}\overline{h}_{y}-\overline{g}_{y}\overline{h}_{x}$ $=0$
$\overline{g}_{y}\overline{h}_{z}-\overline{g}_{z}\overline{h}_{y}$ $=z\,\overline{\varphi}_{y}$
$\overline{g}_{x}\overline{h}_{z}-\overline{g}_{z}\overline{h}_{x}$ $=z\,\overline{\varphi}_{x}$
$\overline{f}_{x}\overline{h}_{y}-\overline{f}_{y}\overline{h}_{x}$ $=0$
$\overline{f}_{y}\overline{h}_{z}-\overline{f}_{z}\overline{h}_{y}$ $=-2z\,\overline{\varphi}_{y}$
$\overline{f}_{x}\overline{h}_{z}-\overline{f}_{z}\overline{h}_{x}$ $=-2z\,\overline{\varphi}_{x}\,.$
By [3, Corollary 2] or [2, Lemma 1], we deduce that $\overline{f},\overline{g},\overline{h}$ are mutually algebraically independent. On the other hand, from the hypothesis $\overline{\varphi}\in F[x,y,z]$ is a homogeneous polynomial of positive degree such that $\overline{\varphi}\not\in F[z]$, it is not difficult to see that none of the polynomials $\overline{f},\overline{g},\overline{h}$ belongs to the subalgebra generated by the other two polynomials. Therefore, by [2, Corollary 8] $(f,g,h)$ is wild. ∎
We give the following definition in order to determine the highest weighted homogeneous component of the polynomial $\varphi=p(xz+y^{2},z)$.
Definition 1.
Let us fix algebraically independent indeterminates $t_{1},\dots,t_{n}$ over $F$. For $k\in\mathbb{Z}^{n}_{\geq 1}$, we write $t^{k}$ to denote the monomial $t_{1}^{k_{1}}\cdots t_{n}^{k_{n}}$. A polynomial $h\in F[t_{1},\dots,t_{n}]$ can be written as $h=\sum_{k}a_{k}t^{k}$.
  1. (1)
    The support of $h$, denoted by $\mathrm{supp}(h)$, is defined as the set
    $\mathrm{supp}(h)=\{k\in\mathbb{Z}^{n}_{\geq 0}:a_{k}\neq 0\}\,.$
  2. (2)
    Let $v=(v_{1},\dots,v_{n})\in\mathbb{Z}^{n}_{\geq 1}$. The degree of $h$ with respect to $v$, denoted by $d_{v}(h)$, is the maximum of the scalar product $\langle v,k\rangle$ such that $a_{k}\neq 0$. We denote by $h^{v}$ the sum of those terms $a_{k}t^{k}$ such that $\langle v,k\rangle=d_{v}(p)$. In particular, if $v=(1,\dots,1)$ then $\deg(h)=d_{v}(h)$ and $h^{v}=\overline{h}$.
  3. (3)
    We say that $h$ is weighted homogeneous with respect to $v$ when $h^{v}=h$.
Lemma 2.
Let $p\in F[t_{1},t_{2}]$ and $v=(2,1)$. If $\varphi=p(xz+y^{2},z)\in F[x,y,z]$, then $\deg(\varphi)=d_{v}(p)$ and $\overline{\varphi}=p^{v}(xz+y^{2},z)$. In particular, if $p$ is weighted homogeneous with respect to $v=(2,1)$, then $\varphi$ is homogeneous.
Proof.
Let us write the polymomial $p$ as $p=\sum_{k}a_{k}t^{k}$. One has
$\varphi=\sum_{k}a_{k}(xz+y^{2})^{k_{1}}z^{k_{2}}=\sum_{k}a_{k}\sum_{n=0}^{k_{1}}\binom{k_{1}}{n}x^{k_{1}-n}y^{2n}z^{k_{1}+k_{2}-n}\,.$
From the last equality, we observe that
(3) $\mathrm{supp}(\varphi)=\Big{\{}(k_{1}-n,2n,k_{1}+k_{2}-n):k\in\mathrm{supp}(p)\,\textnormal{and}\,\,0\leq n\leq k_{1}\Big{\}}\,.$
Therefore,
$\deg(\varphi)$ $=\max\Big{\{}k_{1}-n+2n+k_{1}+k_{2}-n:(k_{1},k_{2})\in\mathrm{supp}(p)\,\textnormal{and}\,0\leq n\leq k_{1}\Big{\}}$
$=\max\Big{\{}2k_{1}+k_{2}:(k_{1},k_{2})\in\mathrm{supp}(p)\Big{\}}$
(4) $=d_{v}(p)\,.$
Hence, by the equalities given in (3) and (4), the highest degree homogeneous component of $\varphi$ is computed as follows
$\overline{\varphi}=$ $\sum_{\begin{subarray}{c}k\in\mathrm{supp}(p)\\ 0\leq n\leq k_{1}\\ 2k_{1}+k_{2}=d_{v}(p)\end{subarray}}a_{k}\sum_{n=0}^{k_{1}}\binom{k_{1}}{n}x^{k_{1}-n}y^{2n}z^{k_{1}+k_{2}-n}$
$=$ $\sum_{\begin{subarray}{c}k\in\mathrm{supp}(p)\\ 0\leq n\leq k_{1}\\ 2k_{1}+k_{2}=d_{v}(p)\end{subarray}}a_{k}\left(\sum_{n=0}^{k_{1}}\binom{k_{1}}{n}x^{k_{1}-n}z^{k_{1}-n}y^{2n}\right)z^{k_{2}}$
$=$ $\sum_{\begin{subarray}{c}k\in\mathrm{supp}(p)\\ 0\leq n\leq k_{1}\\ 2k_{1}+k_{2}=d_{v}(p)\end{subarray}}a_{k}\left(\sum_{n=0}^{k_{1}}\binom{k_{1}}{n}(xz)^{k_{1}-n}y^{2n}\right)z^{k_{2}}$
$=$ $\sum_{\begin{subarray}{c}k\in\mathrm{supp}(p)\\ 2k_{1}+k_{2}=d_{v}(p)\end{subarray}}\left(xz+y^{2}\right)^{k_{1}}z^{k_{2}}$
$=$ $p^{v}(xz+y^{2},z)\,,$
and it proves the assertion. ∎
Example 1.
If $\varphi(x,y,z)=p(xz+y^{2},z)$, then it is not always true that $\overline{\varphi}(x,y,z)=\overline{p}(xz+y^{2},z)$. Indeed, consider $p(t_{1},t_{2})=t_{1}^{2}-t_{2}^{3}+t_{1}t_{2}^{2}\in F[s,t]$, then by a immediate computation one has
$\overline{\varphi}=\left(xz+y^{2}\right)^{2}+\left(xz+y^{2}\right)z^{2}\neq-z^{3}+\left(xz+y^{2}\right)z^{2}=\overline{p}(xz+y^{2},z)\,.$
However, it is easy to verify that if $v=(2,1)$, then $d_{v}(p)=4$ and $p^{v}=t_{1}^{2}+t_{1}t_{2}^{2}$. Moreover, by direct computation one has
$\overline{\varphi}=\left(xz+y^{2}\right)^{2}+\left(xz+y^{2}\right)z^{2}=p^{v}(xz+y^{2},z)\,.$
Corollary 1.
Let $p\in F[t_{1},t_{2}]$ and set $\varphi=p(xz+y^{2},z)$ in $F[x,y,z]$. If the partial derivative $(p^{v})_{t_{1}}$ is nonzero, then the corresponding ring homomorphism $(f,g,h)$ given in (2) is a wild automorphism.
Proof.
By Proposition 2, $(f,g,h)$ is an automorphism and Lemma 2 establishes the equality $\overline{\varphi}=p^{v}(xz+y^{2},z)$. Hence the fact that $(p^{v})_{t_{1}}\neq 0$ implies that $\overline{\varphi}_{x}\neq 0$ and $\overline{\varphi}_{y}\neq 0$. By Proposition 3, the automorphism $(f,g,h)$ is wild. ∎
Example 2.
Let $\varphi=(xz+y^{2})^{n}$ with $n\geq 1$. Corollary 1 implies that
$f(x,y,z)$ $=x-2y(xz+y^{2})^{n}-z(xz+y^{2})^{2n},$
$g(x,y,z)$ $=y+z(xz+y^{2})^{n},$
$h(x,y,z)$ $=z,$
is a wild automorphism.
Example 3.
Let $\varphi(x,y,z)=p(xz+y^{2},z)$, where $p(t_{1},t_{2})=t_{1}^{2}-t_{2}^{3}+t_{1}t_{2}^{2}\in F[s,t]$ as in Example 1. Since $p^{v}=t_{1}^{2}+t_{1}t_{2}^{2}$ and $(p^{v})_{t_{1}}=2t_{1}+t_{1}^{2}\neq 0$, Corollary 1 implies that corresponding ring homomorphism $(f,g,h)$ is a wild automorphism.
3. Statement of the main results
The main results in this work are the following:
Theorem 1.
The linear PDE
(5) $-2y\varphi_{x}+z\varphi_{y}=0$
has as general solution in $F[x,y,z]$, polynomials that can be written as:
$\varphi=p(xz+y^{2},z)\,,$
where $p\in F[t_{1},t_{2}]$.
Theorem 2.
Let $\varphi\in F[x,y,z]$ and set
$f(x,y,z)$ $=x-2y\varphi-z\varphi^{2}\,,$
$g(x,y,z)$ $=y+z\varphi\,,$
$h(x,y,z)$ $=z\,.$
Then, the determinant of the Jacobian matrix $J(f,g,h)$ is a nonzero constant in $F$ if and only if $(f,g,h)$ is an automorphism on $F[x,y,z]$. In this case, there exists a polynomial $p\in F[t_{1},t_{2}]$ such that $\varphi=p(xz+y^{2},z)$, moreover the inverse of $(f,g,h)$ is given in (2).
The proofs of these results are given after the following previous results.
Lemma 3.
Let $\varphi=\varphi^{0}+\varphi^{1}+\cdots+\varphi^{n}$ be a solution of the linear PDE:
(6) $A(x,y,z)\varphi_{x}+B(x,y,z)\varphi_{y}+C(x,y,x)\varphi_{z}=0\,,$
where each $\varphi^{k}\in F[x,y.z]$ is the homogeneous component of $\varphi$ of degree $k$ for $0\leq k\leq n$, and $A,B,C$ are homogeneous polynomials of the same degree. Then each homogeneous component $\varphi^{k}$ is also a solution of the equation (6).
Proof.
Replacing $\varphi=\sum_{k=0}^{n}\varphi^{k}$ into the linear PDE (5), we obtain:
$\sum_{k=0}^{n}\Big{(}A(x,y,z)\varphi^{k}_{x}+B(x,y,z)\varphi^{k}_{y}+C(x,y,x)\varphi^{k}_{z}\Big{)}=0$
which is a sum of homogeneous polynomials of different degrees, therefore:
$A(x,y,z)\varphi^{k}_{x}+B(x,y,z)\varphi^{k}_{y}+C(x,y,x)\varphi^{k}_{z}=0\,\,\,\textnormal{for all}\,\,k=0,1,\dots,n\,.$
By the previous lemma, to find the general solution of the equation (5), it is enough to deal with its homogeneous polynomials components. Let $\varphi\in F[x,y,z]$ be a homogeneous polynomial of degree $n$, if we consider the canonical isomorphism of rings $F[x,y,z]\cong F[x,y][z]$, then we can suppose that $\varphi$ is written as:
(7) $\varphi=\varphi^{n}+\varphi^{n-1}z+\varphi^{n-2}z^{2}+\dots+\varphi^{1}z^{n-1}+\varphi^{0}z^{n}=\sum_{k=0}^{n-1}\varphi^{n-k}z^{k}+\varphi^{0}z^{n}$
with $\varphi^{k}\in F[x,y]$ homogeneous polynomials of degree $k$ for $k=0,1,\dots,n$.
Lemma 4.
Let $\varphi$ be a homogeneous polynomial in $F[x,y,z]$ of degree $n$ that is a solution of the linear PDE (5). Then, $\varphi^{n}=ay^{n}$ and $\varphi^{1}=bx$, where $a$ and $b$ are constants in $F$, and moreover
$2y\varphi^{n-k-1}_{x}=\varphi^{n-k}_{y}$
for all $k=0,1,\dots,n-2$.
Proof.
Let us derivate $\varphi$ with respect to $x$ and $y$:
$\varphi_{x}$ $=\sum_{k=0}^{n-1}\varphi^{n-k}_{x}z^{k}=\varphi_{x}^{n}+\sum_{k=1}^{n-1}\varphi^{n-k}_{x}z^{k}\,,$
$\varphi_{y}$ $=\sum_{k=0}^{n-1}\varphi^{n-k}_{y}z^{k}=\sum_{k=0}^{n-2}\varphi^{n-k}_{y}z^{k}+\varphi^{1}_{y}z^{n-1}\,.$
Substituting these equalities in the equation (5), we have
$-2y\varphi_{x}^{n}+\sum_{k=1}^{n-1}-2y\varphi^{n-k}_{x}z^{k}+\sum_{k=0}^{n-2}\varphi^{n-k}_{y}z^{k+1}+\varphi^{1}_{y}z^{n}=0\,,$
and by index rescaling, we get
$-2y\varphi_{x}^{n}+\sum_{k=0}^{n-2}-2y\varphi^{n-k-1}_{x}z^{k+1}+\sum_{k=0}^{n-2}\varphi^{n-k}_{y}z^{k+1}+\varphi^{1}_{y}z^{n}=0\,.$
Hence,
$-2y\varphi_{x}^{n}+\sum_{k=0}^{n-2}\left(-2y\varphi^{n-k-1}_{x}+\varphi^{n-k}_{y}\right)z^{k+1}+\varphi^{1}_{y}z^{n}=0\,.$
This equation gives the equalities $\varphi^{n}_{x}=0$ and $\varphi^{1}_{y}=0$, so $\varphi^{n}=ay^{n}$ and $\varphi^{1}=bx$ for some constants $a,b\in F$. Moreover,
(8) $2y\varphi^{n-k-1}_{x}=\varphi^{n-k}_{y}\,\,\,\textnormal{para todo}\,\,\,k=0,1,\dots,n-2\,.$
The following result, will enable us to know how a solution can be written of the equation (5).
Proposition 4.
Let $\varphi\in F[x,y,z]$ be a homogeneous solution of the linear PDE (5). We have the following assertions:
  1. ($a$)
    If $\varphi$ has degree $2n$, then
    $\varphi=a(xz+y^{2})^{n}+z\psi$
    where $a\in F$ and $\psi\in F[x,y,z]$ is a homogeneous polynomial which is also a solution of the equation (5) of degree $12n-$ 1.
  2. ($b$)
    If $\varphi$ has degree $12n+$ 1, then
    $\varphi=z\psi$
    where $\psi\in F[x,y,z]$ is homogeneous of degree $2n$ which is also a solution of (5) of degree $2n$.
Proof.
($a$). Let us suppose that:
(9) $\varphi=\varphi^{2n}+\varphi^{2n-1}z+\cdots+\varphi^{1}z^{2n-1}+\varphi^{0}z^{2n}$
with $\varphi^{2n-k}$ be a homogeneous polynomial in $F[x,y]$ of degree $2n-k$ for $0\leq k\leq 2n$. By Lemma 4, we get the following equalities:
(10) $\varphi^{2n}$ $=ay^{2n}\,\,\,\,\textnormal{and}\,\,\,\,\varphi^{1}=bx$
(11) $2y\varphi^{2n-k-1}_{x}$ $=\varphi^{2n-k}_{y}\,\,\,\textnormal{para todo}\,\,\,k=0,1,\dots,2n-2\,.$
We claim that there are constants $a,a^{2n-k}_{j}$, with $1\leq k\leq n$ y $10\leq j\leq k-$ 1, such that:
(12) $\varphi^{2n-k}$ $=a\binom{n}{k}y^{2n-2k}x^{k}+\sum_{j=0}^{k-1}a^{2n-k}_{k-1-j}y^{2n-2k+1+j}x^{k-1-j}$
Indeed, taking $k=0$, from these equalities (10) and (11) we find that:
$2y\varphi^{2n-1}_{x}=\varphi^{2n}_{y}=2nay^{2n-1}$
so $\varphi_{x}^{2n-1}=nay^{2n-2}$. Hence, by integrating with respect to $x$, we obtain
$\varphi^{2n-1}=nay^{2n-2}x+a^{2n-1}_{0}y^{2n-1}\,,$
which can be written as
$\varphi^{2n-1}=a\binom{n}{1}y^{2n-2}x+a^{2n-1}_{0}y^{2n-1}\,,$
thus (12) is true for $k=1$. Suppose we have (12). Taking the derivative with respect to $y$, we have
$\varphi^{2n-k}_{y}=2a\binom{n}{k}(n-k)y^{2n-2k-1}x^{k}+\sum_{j=0}^{k-1}\,.(2n-2k+1+j)a^{2n-k}_{k-1-j}y^{2n-2k+j}x^{k-1-j}$
From the equality (11), we get
$\varphi^{2n-k-1}_{x}=a\binom{n}{k}(n-k)y^{2n-2k-2}x^{k}+\sum_{j=0}^{k-1}\frac{(2n-2k+1+j)a^{2n-k}_{k-1-j}}{2}y^{2n-2k-1+j}x^{k-1-j}\,.$
Integrating with respect to $x$, there is a constant $a^{2n-k-1}_{0}\in F$ such that
$\varphi^{2n-k-1}=a\binom{n}{k+1}y^{2n-2k-2}x^{k+1}+\sum_{j=0}^{k-1}\frac{(2n-2k+1+j)a^{2n-k}_{k-1-j}}{2(k-1-j)}y^{2n-2k-1+j}x^{k-j}+a^{2n-k-1}_{0}y^{2n-k-1}\,.$
Defining $a^{2n-k-1}_{k-j}=\frac{(2n-2k+1+j)a^{2n-k}_{k-1-j}}{2(k-1-j)}$, we get
$\varphi^{2n-k-1}=a\binom{n}{k+1}y^{2n-2k-2}x^{k+1}+\sum_{j=0}^{k}a^{2n-k-1}_{k-j}y^{2n-2k-1+j}x^{k-j}$
as wished, and this completes the recurrence. Now, replacing the expressions obtained in (10) and (12) in the equation (9), we have
$\varphi$ $=\sum_{k=0}^{n}\varphi^{2n-k}z^{k}+\sum_{k=n+1}^{2n}\varphi^{2n-k}z^{k}$
$=\varphi^{2n}+\sum_{k=1}^{n}\varphi^{2n-k}z^{k}+\sum_{k=n+1}^{2n}\varphi^{2n-k}z^{k}$
$=ay^{2n}+\sum_{k=1}^{n}\left(a\binom{n}{k}y^{2n-2k}x^{k}+\sum_{j=0}^{k-1}a^{2n-k}_{k-1-j}y^{2n-2k+1+j}x^{k-1-j}\right)z^{k}+\sum_{k=n+1}^{2n}\varphi^{2n-k}z^{k}$
$=a\sum_{k=0}^{n}\binom{n}{k}y^{2n-2k}x^{k}z^{k}+\sum_{k=1}^{n}\left(\sum_{j=0}^{k-1}a^{2n-k}_{k-1-j}y^{2n-2k+1+j}x^{k-1-j}\right)z^{k}+\sum_{k=n+1}^{2n}\varphi^{2n-k}z^{k}\,.$
From the last equality, we deduce that
$\varphi=a(xz+y^{2})^{n}+z\psi,$
where $\psi\in F[x,y,z]$ is a homogeneous polynomial of degree equal to $12n-$ 1. By the linearity of the solutions of the equation (5), we obtain that $z\psi$ is a solution, hence $\psi$ is also a solution. ($b$) Let us suppose that:
(13) $\varphi=\varphi^{2n+1}+\varphi^{2n}+\cdots+\varphi^{1}z^{2n}+\varphi^{0}z^{2n+1}\,,$
where $\varphi^{2n-k}$ is a homogeneous polynomial in $F[x,y]$ of degree equal to $2n+1-k$, for all $10\leq k\leq 2n+$ 1. By Lemma 4, we get the following equalities:
(14) $\varphi^{2n+1}$ $=ay^{2n+1}$
(15) $2y\varphi^{2n-k}_{x}$ $=\varphi^{2n+1-k}_{y}\,\,\,\textnormal{para todo}\,\,\,k=0,1,\dots,2n-1\,.$
As in the even case, we claim that there are constants $a,a_{j}^{2n-k}$, with $10\leq k\leq n-$ 1 and $0\leq j\leq k$, such that:
(16) $\varphi^{2n-k}=\frac{(2n-2k+1)\cdots(2n-1)(2n+1)}{2^{k}\cdot(k+1)!}ay^{2n-2k-1}x^{k+1}+\sum_{j=0}^{k}b^{2n-k}_{k-j}y^{2n-2k+j}x^{k-j}\,.$
Replacing $k=0$ in the equation (14) and joining with (15), we have
$2y\varphi^{2n}_{x}=\varphi^{2n+1}_{y}=(2n+1)ay^{2n}\,,$
hence $\varphi_{x}^{2n}=(2n+1)ay^{2n-1}$. Integrating with respect to $x$, we obtain
$\varphi^{2n}=(2n+1)ay^{2n-1}x+b^{2n}_{0}y^{2n}$
thus (16) is true for $k=1$. Suppose we have (16). Taking the derivative with respect to $y$, we have
$\varphi^{2n-k}_{y}=\frac{(2n-2k-1)\cdots(2n-1)(2n+1)}{2^{k}\cdot(k+1)!}ay^{2n-2k-2}x^{k+1}+\sum_{j=0}^{k}(2n-2k+j)b^{2n-k}_{k-j}y^{2n-2k+j-1}x^{k-j}\,.$
From the equality (15), we get
$\varphi^{2n-k-1}_{x}=\frac{(2n-2k-1)\cdots(2n-1)(2n+1)}{2^{k+1}\cdot(k+1)!}ay^{2n-2k-3}x^{k+1}+\sum_{j=0}^{k}\frac{(2n-2k+j)b^{2n-k}_{k-j}}{2}y^{2n-2k+j-2}x^{k-j}\,.$
Integrating with respect to $x$, there is a constant $b^{2n-k-1}_{0}\in F$ such that $\varphi^{2n-k-1}$ is equal to
$\frac{(2n-2k-1)\cdots(2n-1)(2n+1)}{2^{k+1}\cdot(k+2)!}ay^{2n-2k-3}x^{k+2}+\sum_{j=0}^{k}\frac{(2n-2k+j)b^{2n-k}_{k-j}}{2(k-j+1)}y^{2n-2k+j-2}x^{k-j+1}+b^{2n-k-1}_{0}y^{2n-k-1}\,.$
Defining $b^{2n-k-1}_{k-j}=\frac{(2n-2k+j)b^{2n-k}_{k-j}}{2(k-j+1)}$, we get
$\varphi^{2n-k-1}=\frac{(2n-2k+-1)\cdots(2n-1)(2n+1)}{2^{k+1}\cdot(k+2)!}ay^{2n-2k-3}x^{k+2}+\sum_{j=0}^{k+1}\frac{(2n-2k+j)b^{2n-k}_{k-j}}{2(k-j+1)}y^{2n-2k+j-2}x^{k-j+1}$
as wished, and this completes the recurrence. In particular, replacing $k=n-1$ in (16), we obtain:
$\varphi^{n+1}=\frac{1\cdot 3\cdots(2n+1)}{2^{n-1}\cdot n!}ayx^{n}+\sum_{j=0}^{n-1}b^{n+1}_{n-1-j}y^{j+2}x^{n-1-j}\,.$
We solve the equation (15) for $k=n$ and we get:
$-2y\varphi^{n}_{x}$ $=\varphi^{n+1}_{y}\,,$
(17) $-2y\varphi^{n}_{x}$ $=\frac{1\cdot 3\cdots(2n+1)}{2^{n-1}\cdot n!}ax^{n}+\sum_{j=0}^{n-1}(j+2)b^{n+1}_{n-1-j}y^{j+1}x^{n-1-j}\,.$
From the equality (17), we deduce that $a=0$ and as consequence from this we obtain that $\varphi^{2n+1}=0$. Therefore,
$\varphi=z\varphi^{2n}+z^{2}\varphi^{2n-1}+\cdots+z^{2n}\varphi^{1}+\varphi^{0}z^{2n+1}=z\psi\,,$
where $\psi\in F[x,y,z]$ is a homogeneous polynomial of degree $2n$, and as in the previous case, we obtain that $\psi$ is also a solution. ∎
4. Proof of the main results
Proof of Theorem 1.
Suppose that $\varphi=p(xz+y^{2},z)$, where $p$ is a polynomial in $F[t_{1},t_{2}]$. Replacing the partial derivatives
$\varphi_{x}$ $=z\cdot p_{s}(xz+y^{2},z)\,,$
$\varphi_{y}$ $=2y\cdot p_{s}(xz+y^{2},z)\,.$
in the linear PDE (5), one verifies that $\varphi=p(xz+y^{2},z)$ is a solution. Reciprocally, suppose that $\varphi\in F[x,y,z]$ is a solution of the equation (5). We shall prove that $\varphi$ has the form $\varphi=p(xz+y^{2},z)$, where $p$ is a polynomial in $F[t_{1},t_{2}]$. By Lemma 3, we can assume that $\varphi\in F[x,y,z]$ is a homogeneous polynomial. More precisely, we shall prove by recurrence the following:
(18) $\varphi=\left\{\begin{array}[]{rcl}\sum_{k=0}^{n}a_{2k}z^{2k}(xz+y^{2})^{2n-2k}&\mbox{if}&\deg\varphi=2n\,,\\ &&\\ \sum_{k=1}^{n}a_{2k-1}z^{2k-1}(xz+y^{2})^{2n-2k}&\mbox{if}&\deg\varphi=2n-1\,.\end{array}\right.$
Indeed, applying Proposition 4 for $n=1$, we obtain:
$\varphi=\left\{\begin{array}[]{ccl}a(xz+y^{2})+z\psi&\mbox{if}&\deg\varphi=2\,,\\ &&\\ z\psi_{1}&\mbox{if}&\deg\varphi=1\,,\end{array}\right.$
where $\psi$ and $\psi_{1}$ are homogeneous polynomials with $\deg\psi=1$ and $\deg\psi_{1}=0$. Then there are constants $b,c\in F$ such that $\psi=cz$ and $\psi_{1}=b\in F$, so (18) is true for $n=1$.
Let us suppose that (18) holds for homogeneous polynomial solutions of degree $2n$ or $12n-$ 1. If $\deg\varphi=2n+2$, then by Proposition 4 ($a$), we obtain that $\varphi=a(xz+y^{2})^{2n+2}+z\psi$, where $\psi$ is a solution of the equation (5). By the inductive hypothesis, we have
$\psi=\sum_{k=1}^{n}a_{2k+1}z^{2k-1}(xz+y^{2})^{2n-2k}\,.$
Hence,
$\varphi$ $=a(xz+y^{2})^{2n+2}+z\psi$
$=a(xz+y^{2})^{2n+2}+z\left(\sum_{k=1}^{n}a_{2k+1}z^{2k-1}(xz+y^{2})^{2n-2k}\right)\,.$
If $\deg\varphi=2n+1$, then by Proposition 4 ($b$), we obtain that $\varphi=z\psi$, where $\psi$ is a solution of the equation (5) with $\deg\psi=2n$. By the inductive hypothesis, one has
$\psi=\sum_{k=0}^{n}a_{2k}z^{2k}(xz+y^{2})^{2n-2k}\,.$
Hence,
$\varphi$ $=z\psi$
$=z\left(\sum_{k=0}^{n}a_{2k}z^{2k}(xz+y^{2})^{2n-2k}\right)$
$=\sum_{k=0}^{n}a_{2k}z^{2k+1}(xz+y^{2})^{2n-2k}$
$=\sum_{k=1}^{n+1}a_{2k-2}z^{2k-1}(xz+y^{2})^{2(n+1)-2k}\,,$
which completes the proof. ∎
Proof of the Theorem 2.
One implication is straightforward. Suppose that $\det J(f,g,h)$ is a nonzero constant. By Proposition 1, $\varphi$ satisfies the linear PDE (5). By Theorem 1, the solution of the equation (5) has the form $\varphi=p(xz+y^{2},z)$ where $p$ is a polynomial in $F[t_{1},t_{2}]$. Finally, by Proposition 2, $(f,g,h)$ is an automorphism and its inverse is given in (2). ∎
5. Conclusions
Theorem 2 says that the automorphisms of the form (2) satisfy the well know Jacobian conjecture via the approach of solving first order partial differential equations. A further problem is whether those automorphisms might be naturally extended on polynomial rings with more than three variables satisfying the aforementioned result.

### v2

@@
 Abstract.
-We define a family of polynomial ring homomorphisms generalising the Nagata automorphism, and relate the condition of non-vanishing of their Jacobian determinant with a certain linear partial differential equation.
-We also give a condition on that family to get wild automorphisms.
+We define a family of polynomial ring homomorphisms generalizing the well-known Nagata automorphism.
+We establish necessary and sufficient conditions under which these homomorphisms are automorphisms, and verify that they satisfy the Jacobian conjecture.
+Additionally, we provide a necessary condition within this family to obtain wild automorphisms, and independently derive a property related to the upper semicontinuity of the Łojasiewicz exponent at infinity.
 Key words and phrases:
-polynomial automorphisms, Nagata automorphism, wild automorphisms, linear PDE
+polynomial automorphisms, Nagata automorphism, wild automorphisms, linear PDE, Łojasiewicz exponent
 2010 Mathematics Subject Classification:
@@
 1. Introduction
-In what follows $F$ will be a field of characteristic 00 unless otherwise stated. The polynomials:
-$f(x,y,z)$ $=x-2y(zx+y^{2})-z(zx+y^{2})^{2},$
-$g(x,y,z)$ $=y+z(zx+y^{2}),$
-$h(x,y,z)$ $=z.$
-define an automorphism on $F[x,y,z]$ called Nagata automorphism [1].
-The determinant of the Jacobian matrix of this automorphism es $1$.
+In what follows $F$ will be a field of characteristic 0 unless otherwise stated. The polynomials
+(1) $\begin{split}f(x,y,z)&=x-2y(zx+y^{2})-z(zx+y^{2})^{2},\\ g(x,y,z)&=y+z(zx+y^{2}),\\ h(x,y,z)&=z.\end{split}$
+define an automorphism of $F[x,y,z]$ called Nagata automorphism [7]. The determinant of the Jacobian matrix of this automorphism turns out to be $1$.
+An automorphism is tame if it is a finite composition of elementary and linear automorphisms. An automorphism that is not tame is called wild [12].
+In 1972, Nagata conjectured that the aforementioned automorphism is wild. This conjecture was proved in [12, Corollary 9].
+In this work, we define a family of ring homomorphisms on $F[x,y,z]$, given in (2), generalizing the Nagata automorphism, which includes the family of automorphisms constructed by Bass [1] in some sense, but is quite different to the family of automorphisms constructed in [6, (1)].
+Our family provides wild automorphisms under some assumptions, see Corollary 1.
+In Theorem 1, we solve a linear partial differential equation, whose solution is related to the non-vanishing of the Jacobian determinant of those automorphisms.
+As a consequence, we obtain a necessary and sufficient condition on homomorphisms of Nagata type with nonzero Jacobian determinant to be automorphisms, which is a particular case of the famous Jacobian conjecture, see Corollary 3.
+Acknowledgments. We are grateful to Carles Bivià-Ausina for valuable comments on an early version of this article.
+2. Nagata polynomial map
 In the following paragraph we recall the notion of tame automorphisms of polynomials.
-Let $1\leq i\leq n$, the $i$th elementary automorphism $\Phi^{(i)}\colon F^{n}\to F^{n}$ is the automorphism defined as
-$\Phi^{(i)}(x_{1},\dots,x_{n})=\Big{(}x_{1},\dots,x_{i-1},x_{i}+f(x_{1},\dots,x_{i-1},x_{i+1},\dots,x_{n}),x_{i+1},\dots,x_{n}\Big{)}$
-and its inverse is the automorphism
-${\Phi^{(i)}}^{-1}(x_{1},\dots,x_{n})=\Big{(}x_{1},\dots,x_{i-1},x_{i}-f(x_{1},\dots,x_{i-1},x_{i+1},\dots,x_{n}),x_{i+1},\dots,x_{n}\Big{)}\,.$
-One may notice that the Jacobian matriz of $\Phi^{(i)}$ has determinant $1$.
-An automorphism is tame if it is a finite composition of elementary and linear automorphisms.
-An automorphism that is not tame is called wild.
-In 1972, Nagata conjectured that the Nagata automorphism is wild.
-In [2, Corollary 9], the authors proved this conjecture.
-In this work, we define ring homomorphisms on $F[x,y,z]$, given in (2), extending the Nagata automorphism which are wild automorphisms under some assumptions, see Corollary 1).
-We also relate the condition of non-vanishing of their Jacobian determinant with a certain linear partial differential equation, see Theorem 1.
-As a consequence, we prove that those ring homomorphisms on $F[x,y,z]$ with nonzero Jacobian determinant, are in fact automorphisms, see Theorem 2.
-2. Nagata polynomial map
+Fix a natural number $n\geq 2$. For each index $1\leq i\leq n$, the $i$th elementary automorphism of $F^{n}$ on itself is the automorphism defined as
+$(x_{1},\dots,x_{n})\mapsto\Big(x_{1},\dots,x_{i-1},x_{i}+f(x_{1},\dots,x_{i-1},x_{i+1},\dots,x_{n}),x_{i+1},\dots,x_{n}\Big)\,$,
+where $f$ is a polynomial map in n1n-1 variables, and its inverse is the automorphism
+$(x_{1},\dots,x_{n})\mapsto\Big(x_{1},\dots,x_{i-1},x_{i}-f(x_{1},\dots,x_{i-1},x_{i+1},\dots,x_{n}),x_{i+1},\dots,x_{n}\Big)\,.$
+It turns out that the Jacobian matrix of each elementary automorphism has determinant $1$.
+Since $F$ is an infinite field, there is a one-to-one correspondence between polynomial maps $F^{n}\to F^{n}$ and $n$-tuples of polynomials in $F[x_{1},\dots,x_{n}]$.
+In what follows, we shall focus on the case $n=3$.
+Definition 1.
 For a polynomial $\varphi\in F[x,y,z]$, we set
 (2) $\begin{split}f(x,y,z)&=x-2y\varphi-z\varphi^{2}\,,\\ g(x,y,z)&=y+z\varphi\,,\\ h(x,y,z)&=z\,.\end{split}$
-The triplet $(f,g,h)$ induces a ring homomorphism on $F[x,y,z]$.
+The triplet $(f,g,h)$ induces an ring homomorphism on $F[x,y,z]$ which will be called Nagata homomorphism.
 Note that $(f,g,h)$ is the identity when one takes $\varphi=0$, so $(f,g,h)$ may be seen as a perturbation of the identity.
-If $\varphi=xz+y^{2}$ then $(f,g,h)$ is the Nagata automorphism.
+If $\varphi=xz+y^{2}$ then $(f,g,h)$ is the Nagata automorphism (1).
+If $\varphi=\lambda(xz+y^{2})$ with $\lambda\in F$, then we obtain a family of automorphisms similar to those constructed by Bass [1], [5, 3.8.1].
 The Milnor number of $(f,g,h)$ is the dimension
@@
 Lemma 1.
-If $(f,g,h)$ is a ring homomorphism as in (2), then $\mu(f,g,h)=1$.
+If $(f,g,h)$ is a Nagata homomorphism, then $\mu(f,g,h)=1$.
 Proof.
@@
 Remark 1.
-Note that $(f,g,h)$ is not always an automorphism for any $\varphi$ despite the fact that $\mu(f,g,h)=1$ by Lemma 1.
-For instance, if $\varphi=x$ or $\varphi=y$, then $(f,g,h)$ is not an automorphism.
-In the first case, $(f,g,h)$ sends both points $(0,0,1)$ and $(-1,1,1)$ to $(0,0,1)$.
-Theorem 2 provides a necessary and sufficient condition on $\varphi$ such that $(f,g,h)$ is an automorphism.
+Note that the Nagata homomorphism $(f,g,h)$ is not always an automorphism for any $\varphi$, despite the fact that $\mu(f,g,h)=1$ by Lemma 1.
+For instance, if $\varphi=x$, then $(f,g,h)$ is not an automorphism, because as a map $(f,g,h)$ sends both points $(0,0,1)$ and (1,1,1)(-1,1,1) to $(0,0,1)$.
+Similarly, $(f,g,h)$ is not an automorphism for $\varphi=y$.
+In contrast, Corollary 3 provides a necessary and sufficient condition on $\varphi$, in order that its corresponding Nagata homomorphism $(f,g,h)$ is an automorphism.
 Proposition 1.
-The Jacobian matrix $J(f,g,h)$ has nonzero determinant in $F$ if and only if $\varphi$ satisfies the linear PDE:
-$-2y\varphi_{x}+z\varphi_{y}=0\,.$
+The Jacobian matrix $J(f,g,h)$ of the Nagata homomorphism $(f,g,h)$, as in (2), has nonzero determinant in $F$ if and only if $\varphi$ satisfies the linear partial differential equation:
+$2y\varphi_{x}+z\varphi_{y}=0\,.$
 Proof.
@@
 $f_{x}$ $=1-2y\varphi_{x}-2z\varphi\varphi_{x},$
-$f_{y}$ $=-2\varphi_{x}-2y\varphi_{y}-2z\varphi\varphi_{y},$
+$f_{y}$ $=-2\varphi-2y\varphi_{y}-2z\varphi\varphi_{y},$
 $g_{x}$ $=z\varphi_{x},$
@@
 Hence the Jacobian matrix $J(f,g,h)$ is
-$\begin{pmatrix}1-2y\varphi_{x}-2z\varphi\varphi_{x}&-2\varphi_{x}-2y\varphi_{y}-2z\varphi\varphi_{y}&f_{z}\\ z\varphi_{x}&1+z\varphi_{y}&g_{z}\\ 0&0&1\end{pmatrix}$
+$\begin{pmatrix}1-2y\varphi_{x}-2z\varphi\varphi_{x}&-2\varphi-2y\varphi_{y}-2z\varphi\varphi_{y}&f_{z}\\ z\varphi_{x}&1+z\varphi_{y}&g_{z}\\ 0&0&1\end{pmatrix}$
 whose determinant is
-$\begin{vmatrix}1-2y\varphi_{x}-2z\varphi\varphi_{x}&-2\varphi_{x}-2y\varphi_{y}-2z\varphi\varphi_{y}\\ z\varphi_{x}&1+z\varphi_{y}\end{vmatrix}=1-2y\varphi_{x}+z\varphi_{y}\,.$
-If $-2y\varphi_{x}+z\varphi_{y}=0$ then the determinant of $J(f,g,h)$ is $1$. Conversely, if the determinant of $J(f,g,h)$ is a constant $c\in F-\{0\}$, then
+$\begin{vmatrix}1-2y\varphi_{x}-2z\varphi\varphi_{x}&-2\varphi-2y\varphi_{y}-2z\varphi\varphi_{y}\\ z\varphi_{x}&1+z\varphi_{y}\end{vmatrix}=1-2y\varphi_{x}+z\varphi_{y}\,.$
+If $2y\varphi_{x}+z\varphi_{y}=0$ then the determinant of $J(f,g,h)$ is $1$. Conversely, if the determinant of $J(f,g,h)$ is a constant $c\in F-\{0\}$, then
 $1-2y\varphi_{x}+z\varphi_{y}=c\,.$
-However this equality is only possible when $c=1$, therefore $-2y\varphi_{x}+z\varphi_{y}=0$. ∎
+However this equality is only possible when $c=1$, therefore $2y\varphi_{x}+z\varphi_{y}=0$. ∎
 Proposition 2.
-Let $p\in F[t_{1},t_{2}]$ and set $\varphi=p(xz+y^{2},z)$ in $F[x,y,z]$. Then the ring homomorphism $(f,g,h)$ given in (2) is an automorphism on $F[x,y,z]$ with inverse:
+Let $p\in F[t_{1},t_{2}]$ and set $\varphi=p(xz+y^{2},z)$ in $F[x,y,z]$. Then the corresponding Nagata homomorphism $(f,g,h)$ is an automorphism of $F[x,y,z]$ with inverse:
 (3) $\begin{split}f^{\prime}(x,y,z)&=x+2y\varphi-z\varphi^{2},\\ g^{\prime}(x,y,z)&=y-z\varphi,\\ h^{\prime}(x,y,z)&=z\,.\end{split}$
@@
 Therefore, the composite $(f,g,h)\circ(f^{\prime},g^{\prime},h^{\prime})$ is the identity. Similarly, one verifies that $(f^{\prime},g^{\prime},h^{\prime})\circ(f,g,h)$ is also the identity. ∎
-For a polynomial $\psi$, we denote by $\overline{\psi}$ the highest degree homogeneous component of $\psi$.
+We give the following definition in order to determine the highest weighted homogeneous component of the polynomial $\varphi=p(xz+y^{2},z)$.
+Definition 2.
+Let us fix algebraically independent indeterminates $t_{1},\dots,t_{n}$ over $F$.
+For $k=(t_{1},\dots,t_{n})\in\mathbb{Z}^{n}_{\geq 1}$, we write $t^{k}$ to denote the monomial $t_{1}^{k_{1}}\cdots t_{n}^{k_{n}}$.
+Any polynomial $h\in F[t_{1},\dots,t_{n}]$ can be written as $h=\sum_{k}a_{k}t^{k}$.
+  1. (1)
+    The support of $h$, denoted by $\mathrm{supp}(h)$, is defined as the set
+    $\mathrm{supp}(h)=\{k\in\mathbb{Z}^{n}_{\geq 0}:a_{k}\neq 0\}\,.$
+  2. (2)
+    Let $v=(v_{1},\dots,v_{n})\in\mathbb{Z}^{n}_{\geq 1}$.
+The degree of $h$ with respect to $v$, denoted by $d_{v}(h)$, is the maximum of the scalar product $\langle v,k\rangle$ such that $a_{k}\neq 0$.
+We denote by $h^{v}$ the sum of those terms $a_{k}t^{k}$ such that $\langle v,k\rangle=d_{v}(h)$.
+In particular, for $v=(1,\dots,1)$, we write
+    $\overline{h}=h^{v}\,$,
+    that is the highest degree homogeneous component of $h$, and in this case we have $d_{v}(h)=\deg(h)$.
+  3. (3)
+    We say that $h$ is weighted homogeneous with respect to $v$ when $h^{v}=h$.
 Proposition 3.
-Let $\varphi\in F[x,y,z]$ and suppose that $(f,g,h)$ is an automorphism on $F[x,y,z]$ of the form (2).
+Let $\varphi\in F[x,y,z]$ and suppose that its corresponding Nagata homomorphism $(f,g,h)$ is an automorphism of $F[x,y,z]$.
 If one the partial derivatives $\overline{\varphi}_{x},\overline{\varphi}_{y}$ is nonzero, then the automorphism $(f,g,h)$ is wild.
@@
 $\overline{f}_{x}\overline{h}_{z}-\overline{f}_{z}\overline{h}_{x}$ $=-2z\,\overline{\varphi}_{x}\,.$
-By [3, Corollary 2] or [2, Lemma 1], we deduce that $\overline{f},\overline{g},\overline{h}$ are mutually algebraically independent.
+By [13, Corollary 2] or [12, Lemma 1], we deduce that $\overline{f},\overline{g},\overline{h}$ are mutually algebraically independent.
 On the other hand, from the hypothesis $\overline{\varphi}\in F[x,y,z]$ is a homogeneous polynomial of positive degree such that $\overline{\varphi}\not\in F[z]$, it is not difficult to see that none of the polynomials $\overline{f},\overline{g},\overline{h}$ belongs to the subalgebra generated by the other two polynomials.
-Therefore, by [2, Corollary 8] $(f,g,h)$ is wild.
+Therefore, by [12, Corollary 8] $(f,g,h)$ is wild.
 ∎
-We give the following definition in order to determine the highest weighted homogeneous component of the polynomial $\varphi=p(xz+y^{2},z)$.
-Definition 1.
-Let us fix algebraically independent indeterminates $t_{1},\dots,t_{n}$ over $F$.
-For $k\in\mathbb{Z}^{n}_{\geq 1}$, we write $t^{k}$ to denote the monomial $t_{1}^{k_{1}}\cdots t_{n}^{k_{n}}$.
-A polynomial $h\in F[t_{1},\dots,t_{n}]$ can be written as $h=\sum_{k}a_{k}t^{k}$.
-  1. (1)
-    The support of $h$, denoted by $\mathrm{supp}(h)$, is defined as the set
-    $\mathrm{supp}(h)=\{k\in\mathbb{Z}^{n}_{\geq 0}:a_{k}\neq 0\}\,.$
-  2. (2)
-    Let $v=(v_{1},\dots,v_{n})\in\mathbb{Z}^{n}_{\geq 1}$.
-The degree of $h$ with respect to $v$, denoted by $d_{v}(h)$, is the maximum of the scalar product $\langle v,k\rangle$ such that $a_{k}\neq 0$.
-We denote by $h^{v}$ the sum of those terms $a_{k}t^{k}$ such that $\langle v,k\rangle=d_{v}(p)$.
-In particular, if $v=(1,\dots,1)$ then $\deg(h)=d_{v}(h)$ and $h^{v}=\overline{h}$.
-  3. (3)
-    We say that $h$ is weighted homogeneous with respect to $v$ when $h^{v}=h$.
 Lemma 2.
@@
 From the last equality, we observe that
-(3) $\mathrm{supp}(\varphi)=\Big{\{}(k_{1}-n,2n,k_{1}+k_{2}-n):k\in\mathrm{supp}(p)\,\textnormal{and}\,\,0\leq n\leq k_{1}\Big{\}}\,.$
+(4) $\mathrm{supp}(\varphi)=\Big\{(k_{1}-n,2n,k_{1}+k_{2}-n):\quad k\in\mathrm{supp}(p)\,\textnormal{and}\,\,0\leq n\leq k_{1}\Big\}\,.$
 Therefore,
-$\deg(\varphi)$ $=\max\Big{\{}k_{1}-n+2n+k_{1}+k_{2}-n:(k_{1},k_{2})\in\mathrm{supp}(p)\,\textnormal{and}\,0\leq n\leq k_{1}\Big{\}}$
-$=\max\Big{\{}2k_{1}+k_{2}:(k_{1},k_{2})\in\mathrm{supp}(p)\Big{\}}$
+$\deg(\varphi)$ $=\max\Big\{k_{1}-n+2n+k_{1}+k_{2}-n:\quad(k_{1},k_{2})\in\mathrm{supp}(p)\,\textnormal{and}\,0\leq n\leq k_{1}\Big\}$
+$=\max\Big\{2k_{1}+k_{2}:\quad(k_{1},k_{2})\in\mathrm{supp}(p)\Big\}$
 (5) $=d_{v}(p)\,.$
@@
 Example 1.
-If $\varphi(x,y,z)=p(xz+y^{2},z)$, then it is not always true that $\overline{\varphi}(x,y,z)=\overline{p}(xz+y^{2},z)$.
-Indeed, consider $p(t_{1},t_{2})=t_{1}^{2}-t_{2}^{3}+t_{1}t_{2}^{2}\in F[s,t]$, then by a immediate computation one has
+If $\varphi(x,y,z)=p(xz+y^{2},z)$, then it is not always true that $\overline{\varphi}$ coincides with $\overline{p}(xz+y^{2},z)$.
+Indeed, consider $p(t_{1},t_{2})=t_{1}^{2}-t_{2}^{3}+t_{1}t_{2}^{2}\in F[t_{1},t_{2}]$, hence $\overline{p}=-t_{2}^{3}+t_{1}t_{2}^{2}$, and by an immediate computation one has
 $\overline{\varphi}=\left(xz+y^{2}\right)^{2}+\left(xz+y^{2}\right)z^{2}\neq-z^{3}+\left(xz+y^{2}\right)z^{2}=\overline{p}(xz+y^{2},z)\,.$
-However, it is easy to verify that if $v=(2,1)$, then $d_{v}(p)=4$ and $p^{v}=t_{1}^{2}+t_{1}t_{2}^{2}$. Moreover, by direct computation one has
-$\overline{\varphi}=\left(xz+y^{2}\right)^{2}+\left(xz+y^{2}\right)z^{2}=p^{v}(xz+y^{2},z)\,.$
+On the other hand, it is easy to verify that if $v=(2,1)$, then $d_{v}(p)=4$ and $p^{v}=t_{1}^{2}+t_{1}t_{2}^{2}$. Finally, by a direct computation one has
+$\overline{\varphi}=p^{v}(xz+y^{2},z)\,.$
 Corollary 1.
 Let $p\in F[t_{1},t_{2}]$ and set $\varphi=p(xz+y^{2},z)$ in $F[x,y,z]$.
-If the partial derivative $(p^{v})_{t_{1}}$ is nonzero, then the corresponding ring homomorphism $(f,g,h)$ given in (2) is a wild automorphism.
+If the partial derivative $(p^{v})_{t_{1}}$ is nonzero, where $v=(2,1)$, then its corresponding Nagata homomorphism $(f,g,h)$ is a wild automorphism.
 Proof.
@@
 Example 3.
-Let $\varphi(x,y,z)=p(xz+y^{2},z)$, where $p(t_{1},t_{2})=t_{1}^{2}-t_{2}^{3}+t_{1}t_{2}^{2}\in F[s,t]$ as in Example 1.
-Since $p^{v}=t_{1}^{2}+t_{1}t_{2}^{2}$ and $(p^{v})_{t_{1}}=2t_{1}+t_{1}^{2}\neq 0$, Corollary 1 implies that corresponding ring homomorphism $(f,g,h)$ is a wild automorphism.
+Let $\varphi(x,y,z)=p(xz+y^{2},z)$, where $p(t_{1},t_{2})=t_{1}^{2}-t_{2}^{3}+t_{1}t_{2}^{2}\in F[t_{1},t_{2}]$ as in Example 1.
+Since $p^{v}=t_{1}^{2}+t_{1}t_{2}^{2}$ and $(p^{v})_{t_{1}}=2t_{1}+t_{1}^{2}\neq 0$, Corollary 1 implies that corresponding Nagata homomorphism $(f,g,h)$ is a wild automorphism.
+Łojasiewicz exponent.— Let $K$ be a normed field.
+For $z=(z_{1},z_{2},\dots,z_{m})\in K^{m}$, we write $|z|=\max_{1\leq i\leq m}|z_{i}|$.
+If $F=(F_{1},F_{2},\dots,F_{n})\colon K^{m}\to K^{n}$ is a polynomial map, then we set $\deg F=\max_{1\leq i\leq n}\deg F_{i}$.
+The Łojasiewicz exponent at infinity of a polynomial map $F\colon K^{m}\to K^{n}$, denoted by $\mathcal{L}_{\infty}(F)$, is defined as the supremum of the real numbers $\alpha$ such that there exist two positive real numbers $C$ and $R$ so that
+$|F(z)|\geq C|z|^{\alpha}$
+for all $|z|\geq R$ in $K^{m}$. One can follow the argument provided in [11, (1.1)] to deduce that
+(6) $\mathcal{L}_{\infty}(F)=\dfrac{1}{\deg(F^{-1})}$
+for every polynomial automorphism $F$.
+For $K=\mathbb{R}$ or $K=\mathbb{C}$, it is well known that $\mathcal{L}_{\infty}(F)$ exists if and only if $F^{-1}(0)$ is compact, and in this case, $\mathcal{L}_{\infty}(F)$ is a rational number [8, Proposition 2.6].
+The computation of this number is generally considered a highly nontrivial problem [4, 9, 11, 2, 3].
+A polynomial deformation of a polynomial $p(t_{1},\dots,t_{m})\in K[t_{1},\dots,t_{m}]$ is a polynomial $H(t_{1},\dots,t_{m},s_{1},\dots,s_{n})\in K[t_{1},\dots,t_{m},s_{1},\dots,s_{n}]$ such that
+$H(t_{1},\dots,t_{m},0,\dots,0)=p(t_{1},\dots,t_{m})\,.$
+For simplicity, we write $s=(s_{1},\dots,s_{n})$, so that the polynomial ring $K[t_{1},\dots,t_{m},s_{1},\dots,s_{n}]$ shall be shortly written as $K[t_{1},\dots,t_{m},s]$, and this allows to set
+$p_{s}(t_{1},\dots,t_{m})=H(t_{1},\dots,t_{m},s_{1},\dots,s_{n})\,.$
+The following is motivated by the lower semicontinuous property of the Łojasiewicz exponent proven for finite holomorphic germs with constant Milnor number [10, Theorem 2.1].
+However, for polynomial maps, we expect the upper semicontinuous property under certain conditions, in view of the fact that Nagata homomorphisms have a constant Milnor number, see Lemma 1.
+Corollary 2.
+Let $p_{s}(t_{1},t_{2})\in K[t_{1},t_{2},s]$ be a polynomial deformation of a polynomial $p(t_{1},t_{2})\in K[t_{1},t_{2}]$.
+Let $\varphi=p(xz+y^{2},z)$, $\varphi_{s}=p_{s}(xz+y^{2},z)$, and suppose that $(f,g,h)$ and $(f_{s},g_{s},h_{s})$ are their corresponding Nagata automorphism.
+Then
+$\mathcal{L}_{\infty}(f_{s},g_{s},h_{s})\leq\mathcal{L}_{\infty}(f,g,h)$
+Proof.
+Since $\mathrm{supp}(p)\subseteq\mathrm{supp}(p_{s})$ with respect to the variables $t_{1}$ and $t_{2}$, we have $\deg(p)\leq\deg(p_{s})$, and for $v=(2,1)$ one has
+$\deg\varphi=d_{v}(p)\leq d_{v}(p_{s})=\deg\varphi_{s}\,.$
+Hence, in view of (6) we have
+$\mathcal{L}_{\infty}(f_{s},g_{s},h_{s})$ $=\dfrac{1}{\deg((f_{s},g_{s},h_{s})^{-1})}$
+$=\dfrac{1}{2\deg\varphi_{s}+1}$
+$\leq\dfrac{1}{2\deg\varphi+1}$
+$=\dfrac{1}{\deg((f,g,h)^{-1})}$
+$=\mathcal{L}_{\infty}(f,g,h)\,.$
 3. Statement of the main results
@@
 Theorem 1.
-The linear PDE
-(5) $-2y\varphi_{x}+z\varphi_{y}=0$
-has as general solution in $F[x,y,z]$, polynomials that can be written as:
-$\varphi=p(xz+y^{2},z)\,,$
-where $p\in F[t_{1},t_{2}]$.
-Theorem 2.
-Let $\varphi\in F[x,y,z]$ and set
-$f(x,y,z)$ $=x-2y\varphi-z\varphi^{2}\,,$
-$g(x,y,z)$ $=y+z\varphi\,,$
-$h(x,y,z)$ $=z\,.$
-Then, the determinant of the Jacobian matrix $J(f,g,h)$ is a nonzero constant in $F$ if and only if $(f,g,h)$ is an automorphism on $F[x,y,z]$.
+The general solution in $F[x,y,z]$ of the linear partial differential equation:
+(7) $2y\varphi_{x}+z\varphi_{y}=0$
+has the form $\varphi=p(xz+y^{2},z)$, where $p\in F[t_{1},t_{2}]$.
+Corollary 3.
+Let $\varphi\in F[x,y,z]$ and let $(f,g,h)$ be its corresponding Nagata homomorphism as in (2).
+Then, the determinant of the Jacobian matrix $J(f,g,h)$ is a nonzero constant in $F$ if and only if $(f,g,h)$ is an automorphism of $F[x,y,z]$.
 In this case, there exists a polynomial $p\in F[t_{1},t_{2}]$ such that $\varphi=p(xz+y^{2},z)$, moreover the inverse of $(f,g,h)$ is given in (3).
-The proofs of these results are given after the following previous results.
+The proofs of these results are given in Section 4. First of all, we shall provide some preliminary results in the next paragraphs.
 Lemma 3.
-Let $\varphi=\varphi^{0}+\varphi^{1}+\cdots+\varphi^{n}$ be a solution of the linear PDE:
-(6) $A(x,y,z)\varphi_{x}+B(x,y,z)\varphi_{y}+C(x,y,x)\varphi_{z}=0\,,$
-where each $\varphi^{k}\in F[x,y.z]$ is the homogeneous component of $\varphi$ of degree $k$ for $0\leq k\leq n$, and $A,B,C$ are homogeneous polynomials of the same degree.
+Let $\varphi=\varphi^{0}+\varphi^{1}+\cdots+\varphi^{n}$ be a solution of the linear partial differential equation:
+(8) $A(x,y,z)\varphi_{x}+B(x,y,z)\varphi_{y}+C(x,y,z)\varphi_{z}=0\,$,
+where each $\varphi^{k}\in F[x,y,z]$ is the homogeneous component of $\varphi$ of degree $k$ for $0\leq k\leq n$, and $A,B,C$ are homogeneous polynomials of the same degree.
 Then each homogeneous component $\varphi^{k}$ is also a solution of the equation (8).
@@
 Replacing $\varphi=\sum_{k=0}^{n}\varphi^{k}$ into the linear PDE (7), we obtain:
-$\sum_{k=0}^{n}\Big{(}A(x,y,z)\varphi^{k}_{x}+B(x,y,z)\varphi^{k}_{y}+C(x,y,x)\varphi^{k}_{z}\Big{)}=0$
-which is a sum of homogeneous polynomials of different degrees, therefore:
-$A(x,y,z)\varphi^{k}_{x}+B(x,y,z)\varphi^{k}_{y}+C(x,y,x)\varphi^{k}_{z}=0\,\,\,\textnormal{for all}\,\,k=0,1,\dots,n\,.$
+$\sum_{k=0}^{n}\Big(A(x,y,z)\varphi^{k}_{x}+B(x,y,z)\varphi^{k}_{y}+C(x,y,x)\varphi^{k}_{z}\Big)=0\,$,
+which is a sum of homogeneous polynomials of different degrees. Therefore,
+$A(x,y,z)\varphi^{k}_{x}+B(x,y,z)\varphi^{k}_{y}+C(x,y,x)\varphi^{k}_{z}=0\,\,\,\textnormal{for all}\,\,k=0,1,\dots,n\,$,
+as wanted. ∎
 By the previous lemma, to find the general solution of the equation (7), it is enough to deal with its homogeneous polynomials components.
-Let $\varphi\in F[x,y,z]$ be a homogeneous polynomial of degree $n$, if we consider the canonical isomorphism of rings $F[x,y,z]\cong F[x,y][z]$, then we can suppose that $\varphi$ is written as:
-(7) $\varphi=\varphi^{n}+\varphi^{n-1}z+\varphi^{n-2}z^{2}+\dots+\varphi^{1}z^{n-1}+\varphi^{0}z^{n}=\sum_{k=0}^{n-1}\varphi^{n-k}z^{k}+\varphi^{0}z^{n}$
-with $\varphi^{k}\in F[x,y]$ homogeneous polynomials of degree $k$ for $k=0,1,\dots,n$.
+Let $\varphi\in F[x,y,z]$ be a homogeneous polynomial of degree $n$, if we consider the canonical isomorphism of rings $F[x,y,z]\cong F[x,y][z]$, then we can suppose that $\varphi$ has the form:
+(9) $\varphi=\phi^{n}+\phi^{n-1}z+\phi^{n-2}z^{2}+\dots+\phi^{1}z^{n-1}+\phi^{0}z^{n}=\sum_{k=0}^{n-1}\phi^{n-k}z^{k}+\phi^{0}z^{n}\,$,
+with $\phi^{k}\in F[x,y]$ homogeneous polynomials of degree $k$ for $k=0,1,\dots,n$.
 Lemma 4.
 Let $\varphi$ be a homogeneous polynomial in $F[x,y,z]$ of degree $n$ that is a solution of the linear PDE (7).
-Then, $\varphi^{n}=ay^{n}$ and $\varphi^{1}=bx$, where $a$ and $b$ are constants in $F$, and moreover
-$2y\varphi^{n-k-1}_{x}=\varphi^{n-k}_{y}$
+Then, $\phi^{n}=ay^{n}$ and $\phi^{1}=bx$, where $a$ and $b$ are constants in $F$, and moreover
+$2y\phi^{n-k-1}_{x}=\phi^{n-k}_{y}$
 for all $k=0,1,\dots,n-2$.
@@
 Let us derivate $\varphi$ with respect to $x$ and $y$:
-$\varphi_{x}$ $=\sum_{k=0}^{n-1}\varphi^{n-k}_{x}z^{k}=\varphi_{x}^{n}+\sum_{k=1}^{n-1}\varphi^{n-k}_{x}z^{k}\,,$
-$\varphi_{y}$ $=\sum_{k=0}^{n-1}\varphi^{n-k}_{y}z^{k}=\sum_{k=0}^{n-2}\varphi^{n-k}_{y}z^{k}+\varphi^{1}_{y}z^{n-1}\,.$
+$\varphi_{x}$ $=\sum_{k=0}^{n-1}\phi^{n-k}_{x}z^{k}=\phi_{x}^{n}+\sum_{k=1}^{n-1}\phi^{n-k}_{x}z^{k}\,$,
+$\varphi_{y}$ $=\sum_{k=0}^{n-1}\phi^{n-k}_{y}z^{k}=\sum_{k=0}^{n-2}\phi^{n-k}_{y}z^{k}+\phi^{1}_{y}z^{n-1}\,.$
 Substituting these equalities in the equation (7), we have
-$-2y\varphi_{x}^{n}+\sum_{k=1}^{n-1}-2y\varphi^{n-k}_{x}z^{k}+\sum_{k=0}^{n-2}\varphi^{n-k}_{y}z^{k+1}+\varphi^{1}_{y}z^{n}=0\,,$
+$2y\phi_{x}^{n}+\sum_{k=1}^{n-1}-2y\phi^{n-k}_{x}z^{k}+\sum_{k=0}^{n-2}\phi^{n-k}_{y}z^{k+1}+\phi^{1}_{y}z^{n}=0\,$,
 and by index rescaling, we get
-$-2y\varphi_{x}^{n}+\sum_{k=0}^{n-2}-2y\varphi^{n-k-1}_{x}z^{k+1}+\sum_{k=0}^{n-2}\varphi^{n-k}_{y}z^{k+1}+\varphi^{1}_{y}z^{n}=0\,.$
+$2y\phi_{x}^{n}+\sum_{k=0}^{n-2}-2y\phi^{n-k-1}_{x}z^{k+1}+\sum_{k=0}^{n-2}\phi^{n-k}_{y}z^{k+1}+\phi^{1}_{y}z^{n}=0\,.$
 Hence,
-$-2y\varphi_{x}^{n}+\sum_{k=0}^{n-2}\left(-2y\varphi^{n-k-1}_{x}+\varphi^{n-k}_{y}\right)z^{k+1}+\varphi^{1}_{y}z^{n}=0\,.$
-This equation gives the equalities $\varphi^{n}_{x}=0$ and $\varphi^{1}_{y}=0$, so $\varphi^{n}=ay^{n}$ and $\varphi^{1}=bx$ for some constants $a,b\in F$. Moreover,
-(8) $2y\varphi^{n-k-1}_{x}=\varphi^{n-k}_{y}\,\,\,\textnormal{para todo}\,\,\,k=0,1,\dots,n-2\,.$
+$-2y\phi_{x}^{n}+\sum_{k=0}^{n-2}\left(-2y\phi^{n-k-1}_{x}+\phi^{n-k}_{y}\right)z^{k+1}+\phi^{1}_{y}z^{n}=0\,.$
+This equation gives the equalities $\phi^{n}_{x}=0$ and $\phi^{1}_{y}=0$, so $\phi^{n}=ay^{n}$ and $\phi^{1}=bx$ for some constants $a,b\in F$. Moreover,
+(10) $2y\phi^{n-k-1}_{x}=\phi^{n-k}_{y}\,\,\,\textnormal{para todo}\,\,\,k=0,1,\dots,n-2\,.$
 The following result, will enable us to know how a solution can be written of the equation (7).
@@
   1. ($a$)
-    If $\varphi$ has degree $2n$, then
-    $\varphi=a(xz+y^{2})^{n}+z\psi$
-    where $a\in F$ and $\psi\in F[x,y,z]$ is a homogeneous polynomial which is also a solution of the equation (5) of degree $12n-$ 1.
+    If $\varphi$ has degree $2n$, then it has the form:
+    $\varphi=a(xz+y^{2})^{n}+z\psi\,$,
+    where $a\in F$ and $\psi\in F[x,y,z]$ is a homogeneous polynomial which is also a solution of the equation (7) of degree 2n12n-1.
   2. ($b$)
-    If $\varphi$ has degree $12n+$ 1, then
-    $\varphi=z\psi$
+    If $\varphi$ has degree $2n+1$, then it has the form:
+    $\varphi=z\psi\,$,
     where $\psi\in F[x,y,z]$ is homogeneous of degree $2n$ which is also a solution of (7) of degree $2n$.
@@
 ($a$). Let us suppose that:
-(9) $\varphi=\varphi^{2n}+\varphi^{2n-1}z+\cdots+\varphi^{1}z^{2n-1}+\varphi^{0}z^{2n}$
-with $\varphi^{2n-k}$ be a homogeneous polynomial in $F[x,y]$ of degree $2n-k$ for $0\leq k\leq 2n$. By Lemma 4, we get the following equalities:
-(10) $\varphi^{2n}$ $=ay^{2n}\,\,\,\,\textnormal{and}\,\,\,\,\varphi^{1}=bx$
-(11) $2y\varphi^{2n-k-1}_{x}$ $=\varphi^{2n-k}_{y}\,\,\,\textnormal{para todo}\,\,\,k=0,1,\dots,2n-2\,.$
-We claim that there are constants $a,a^{2n-k}_{j}$, with $1\leq k\leq n$ y $10\leq j\leq k-$ 1, such that:
-(12) $\varphi^{2n-k}$ $=a\binom{n}{k}y^{2n-2k}x^{k}+\sum_{j=0}^{k-1}a^{2n-k}_{k-1-j}y^{2n-2k+1+j}x^{k-1-j}$
-Indeed, taking $k=0$, from these equalities (10) and (11) we find that:
-$2y\varphi^{2n-1}_{x}=\varphi^{2n}_{y}=2nay^{2n-1}$
-so $\varphi_{x}^{2n-1}=nay^{2n-2}$. Hence, by integrating with respect to $x$, we obtain
-$\varphi^{2n-1}=nay^{2n-2}x+a^{2n-1}_{0}y^{2n-1}\,,$
+(11) $\varphi=\phi^{2n}+\phi^{2n-1}z+\cdots+\phi^{1}z^{2n-1}+\phi^{0}z^{2n}\,$,
+where $\phi^{2n-k}\in F[x,y]$ is a homogeneous polynomial of degree 2nk2n-k for $0\leq k\leq 2n$. By Lemma 4, we get the following equalities:
+(12) $\phi^{2n}$ $=ay^{2n}\,\,\,\,\textnormal{and}\,\,\,\,\phi^{1}=bx$
+(13) $2y\phi^{2n-k-1}_{x}$ $=\phi^{2n-k}_{y}\,\,\,\textnormal{para todo}\,\,\,k=0,1,\dots,2n-2\,.$
+We claim that there are constants $a,a^{2n-k}_{k-1-j}$, with $1\leq k\leq n$ y $0\leq j\leq k-1$, such that:
+(14) $\phi^{2n-k}$ $=a\binom{n}{k}y^{2n-2k}x^{k}+\sum_{j=0}^{k-1}a^{2n-k}_{k-1-j}y^{2n-2k+1+j}x^{k-1-j}$
+Indeed, taking $k=0$, from the equalities (12) and (13), we find that:
+$2y\phi^{2n-1}_{x}=\phi^{2n}_{y}=2nay^{2n-1}\,$,
+so that $\phi_{x}^{2n-1}=nay^{2n-2}$. Hence, integrating with respect to $x$, we obtain
+$\phi^{2n-1}=nay^{2n-2}x+a^{2n-1}_{0}y^{2n-1}\,$,
 which can be written as
-$\varphi^{2n-1}=a\binom{n}{1}y^{2n-2}x+a^{2n-1}_{0}y^{2n-1}\,,$
-thus (12) is true for $k=1$. Suppose we have (12). Taking the derivative with respect to $y$, we have
-$\varphi^{2n-k}_{y}=2a\binom{n}{k}(n-k)y^{2n-2k-1}x^{k}+\sum_{j=0}^{k-1}\,.(2n-2k+1+j)a^{2n-k}_{k-1-j}y^{2n-2k+j}x^{k-1-j}$
+$\phi^{2n-1}=a\binom{n}{1}y^{2n-2}x+a^{2n-1}_{0}y^{2n-1}\,$,
+thus (14) is true for $k=1$. Now, suppose that we have (14). Taking the derivative with respect to $y$, we have
+$\phi^{2n-k}_{y}=2a\binom{n}{k}(n-k)y^{2n-2k-1}x^{k}+\sum_{j=0}^{k-1}\,(2n-2k+1+j)a^{2n-k}_{k-1-j}y^{2n-2k+j}x^{k-1-j}$
+By the equality (13), we get
+$\phi^{2n-k-1}_{x}=a\binom{n}{k}(n-k)y^{2n-2k-2}x^{k}+\sum_{j=0}^{k-1}\frac{(2n-2k+1+j)a^{2n-k}_{k-1-j}}{2}y^{2n-2k-1+j}x^{k-1-j}\,.$
+Integrating with respect to $x$, there exists a constant $a^{2n-k-1}_{0}\in F$ such that
+$\phi^{2n-k-1}=a\binom{n}{k+1}y^{2n-2k-2}x^{k+1}+\sum_{j=0}^{k-1}a^{2n-k-1}_{k-j}y^{2n-2k-1+j}x^{k-j}+a^{2n-k-1}_{0}y^{2n-k-1}\,$,
+where $a^{2n-k-1}_{k-j}=\frac{(2n-2k+1+j)a^{2n-k}_{k-1-j}}{2(k-j)}$ for all $0\leq j\leq k-1$. Thus, we obtain
+$\phi^{2n-k-1}=a\binom{n}{k+1}y^{2n-2k-2}x^{k+1}+\sum_{j=0}^{k}a^{2n-k-1}_{k-j}y^{2n-2k-1+j}x^{k-j}\,$,
+as wished, and this completes the recurrence.
+Now, replacing the expressions (12) and (14) in the equation (11), we have
+$\varphi$ $=\phi^{2n}+\sum_{k=1}^{n}\phi^{2n-k}z^{k}+\sum_{k=n+1}^{2n}\phi^{2n-k}z^{k}$
+$=ay^{2n}+\sum_{k=1}^{n}\left(a\binom{n}{k}y^{2n-2k}x^{k}+\sum_{j=0}^{k-1}a^{2n-k}_{k-1-j}y^{2n-2k+1+j}x^{k-1-j}\right)z^{k}+\sum_{k=n+1}^{2n}\phi^{2n-k}z^{k}$
+$=a\sum_{k=0}^{n}\binom{n}{k}y^{2n-2k}x^{k}z^{k}+\sum_{k=1}^{n}\left(\sum_{j=0}^{k-1}a^{2n-k}_{k-1-j}y^{2n-2k+1+j}x^{k-1-j}\right)z^{k}+\sum_{k=n+1}^{2n}\phi^{2n-k}z^{k}\,.$
+From the last equality, we can write
+$\varphi=a(xz+y^{2})^{n}+z\psi,$
+where $\psi\in F[x,y,z]$ is a homogeneous polynomial of degree equal to 2n12n-1.
+By the linearity of the solutions of the equation (7), we deduce that $z\psi$ is a solution, hence $\psi$ is also a solution.
+($b$) Let us suppose that:
+(15) $\varphi=\phi^{2n+1}+\phi^{2n}z+\cdots+\phi^{1}z^{2n}+\phi^{0}z^{2n+1}\,$,
+where $\phi^{2n+1-k}$ is a homogeneous polynomial in $F[x,y]$ of degree equal to 2n+1k2n+1-k, for all $0\leq k\leq 2n+1$. By Lemma 4, we get the following equalities:
+(16) $\phi^{2n+1}$ $=ay^{2n+1}$
+(17) $2y\phi^{2n-k}_{x}$ $=\phi^{2n+1-k}_{y}\,\,\,\textnormal{para todo}\,\,\,k=0,1,\dots,2n-1\,.$
+As in the even case, we claim that there are constants $a,b_{k-j}^{2n-k}$, with $0\leq k\leq n-1$ and $0\leq j\leq k$, such that:
+(18) $\phi^{2n-k}=P^{2n+1}_{k}ay^{2n-2k-1}x^{k+1}+\sum_{j=0}^{k}b^{2n-k}_{k-j}y^{2n-2k+j}x^{k-j}\,$,
+where $P^{2n+1}_{k}=\frac{(2n+1)\cdot(2n+1-2)\cdots(2n+1-2k)}{2^{k+1}\cdot(k+1)!}$ for all index $0\leq k\leq n-1$.
+Replacing $k=0$ in the equation (16) and joining with (17), we have
+$2y\phi^{2n}_{x}=\phi^{2n+1}_{y}=(2n+1)ay^{2n}\,$,
+hence $\phi_{x}^{2n}=\cfrac{(2n+1)}{2}\,\,ay^{2n-1}$. Integrating with respect to $x$, we obtain
+$\phi^{2n}=\cfrac{(2n+1)}{2}\,ay^{2n-1}x+b^{2n}_{0}y^{2n}$
+thus (18) is true for $k=0$. Suppose that we have (18). Taking the derivative with respect to $y$, we have
+$\phi^{2n-k}_{y}=P^{2n+1}_{k}(2n-2k-1)ay^{2n-2k-2}x^{k+1}+\sum_{j=0}^{k}(2n-2k+j)b^{2n-k}_{k-j}y^{2n-2k+j-1}x^{k-j}\,.$
 From the equality (17), we get
-$\varphi^{2n-k-1}_{x}=a\binom{n}{k}(n-k)y^{2n-2k-2}x^{k}+\sum_{j=0}^{k-1}\frac{(2n-2k+1+j)a^{2n-k}_{k-1-j}}{2}y^{2n-2k-1+j}x^{k-1-j}\,.$
-Integrating with respect to $x$, there is a constant $a^{2n-k-1}_{0}\in F$ such that
-$\varphi^{2n-k-1}=a\binom{n}{k+1}y^{2n-2k-2}x^{k+1}+\sum_{j=0}^{k-1}\frac{(2n-2k+1+j)a^{2n-k}_{k-1-j}}{2(k-1-j)}y^{2n-2k-1+j}x^{k-j}+a^{2n-k-1}_{0}y^{2n-k-1}\,.$
-Defining $a^{2n-k-1}_{k-j}=\frac{(2n-2k+1+j)a^{2n-k}_{k-1-j}}{2(k-1-j)}$, we get
-$\varphi^{2n-k-1}=a\binom{n}{k+1}y^{2n-2k-2}x^{k+1}+\sum_{j=0}^{k}a^{2n-k-1}_{k-j}y^{2n-2k-1+j}x^{k-j}$
-as wished, and this completes the recurrence. Now, replacing the expressions obtained in (10) and (12) in the equation (9), we have
-$\varphi$ $=\sum_{k=0}^{n}\varphi^{2n-k}z^{k}+\sum_{k=n+1}^{2n}\varphi^{2n-k}z^{k}$
-$=\varphi^{2n}+\sum_{k=1}^{n}\varphi^{2n-k}z^{k}+\sum_{k=n+1}^{2n}\varphi^{2n-k}z^{k}$
-$=ay^{2n}+\sum_{k=1}^{n}\left(a\binom{n}{k}y^{2n-2k}x^{k}+\sum_{j=0}^{k-1}a^{2n-k}_{k-1-j}y^{2n-2k+1+j}x^{k-1-j}\right)z^{k}+\sum_{k=n+1}^{2n}\varphi^{2n-k}z^{k}$
-$=a\sum_{k=0}^{n}\binom{n}{k}y^{2n-2k}x^{k}z^{k}+\sum_{k=1}^{n}\left(\sum_{j=0}^{k-1}a^{2n-k}_{k-1-j}y^{2n-2k+1+j}x^{k-1-j}\right)z^{k}+\sum_{k=n+1}^{2n}\varphi^{2n-k}z^{k}\,.$
-From the last equality, we deduce that
-$\varphi=a(xz+y^{2})^{n}+z\psi,$
-where $\psi\in F[x,y,z]$ is a homogeneous polynomial of degree equal to $12n-$ 1.
-By the linearity of the solutions of the equation (5), we obtain that $z\psi$ is a solution, hence $\psi$ is also a solution.
-($b$) Let us suppose that:
-(13) $\varphi=\varphi^{2n+1}+\varphi^{2n}+\cdots+\varphi^{1}z^{2n}+\varphi^{0}z^{2n+1}\,,$
-where $\varphi^{2n-k}$ is a homogeneous polynomial in $F[x,y]$ of degree equal to $2n+1-k$, for all $10\leq k\leq 2n+$ 1. By Lemma 4, we get the following equalities:
-(14) $\varphi^{2n+1}$ $=ay^{2n+1}$
-(15) $2y\varphi^{2n-k}_{x}$ $=\varphi^{2n+1-k}_{y}\,\,\,\textnormal{para todo}\,\,\,k=0,1,\dots,2n-1\,.$
-As in the even case, we claim that there are constants $a,a_{j}^{2n-k}$, with $10\leq k\leq n-$ 1 and $0\leq j\leq k$, such that:
-(16) $\varphi^{2n-k}=\frac{(2n-2k+1)\cdots(2n-1)(2n+1)}{2^{k}\cdot(k+1)!}ay^{2n-2k-1}x^{k+1}+\sum_{j=0}^{k}b^{2n-k}_{k-j}y^{2n-2k+j}x^{k-j}\,.$
-Replacing $k=0$ in the equation (14) and joining with (15), we have
-$2y\varphi^{2n}_{x}=\varphi^{2n+1}_{y}=(2n+1)ay^{2n}\,,$
-hence $\varphi_{x}^{2n}=(2n+1)ay^{2n-1}$. Integrating with respect to $x$, we obtain
-$\varphi^{2n}=(2n+1)ay^{2n-1}x+b^{2n}_{0}y^{2n}$
-thus (16) is true for $k=1$. Suppose we have (16). Taking the derivative with respect to $y$, we have
-$\varphi^{2n-k}_{y}=\frac{(2n-2k-1)\cdots(2n-1)(2n+1)}{2^{k}\cdot(k+1)!}ay^{2n-2k-2}x^{k+1}+\sum_{j=0}^{k}(2n-2k+j)b^{2n-k}_{k-j}y^{2n-2k+j-1}x^{k-j}\,.$
-From the equality (15), we get
-$\varphi^{2n-k-1}_{x}=\frac{(2n-2k-1)\cdots(2n-1)(2n+1)}{2^{k+1}\cdot(k+1)!}ay^{2n-2k-3}x^{k+1}+\sum_{j=0}^{k}\frac{(2n-2k+j)b^{2n-k}_{k-j}}{2}y^{2n-2k+j-2}x^{k-j}\,.$
-Integrating with respect to $x$, there is a constant $b^{2n-k-1}_{0}\in F$ such that $\varphi^{2n-k-1}$ is equal to
-$\frac{(2n-2k-1)\cdots(2n-1)(2n+1)}{2^{k+1}\cdot(k+2)!}ay^{2n-2k-3}x^{k+2}+\sum_{j=0}^{k}\frac{(2n-2k+j)b^{2n-k}_{k-j}}{2(k-j+1)}y^{2n-2k+j-2}x^{k-j+1}+b^{2n-k-1}_{0}y^{2n-k-1}\,.$
-Defining $b^{2n-k-1}_{k-j}=\frac{(2n-2k+j)b^{2n-k}_{k-j}}{2(k-j+1)}$, we get
-$\varphi^{2n-k-1}=\frac{(2n-2k+-1)\cdots(2n-1)(2n+1)}{2^{k+1}\cdot(k+2)!}ay^{2n-2k-3}x^{k+2}+\sum_{j=0}^{k+1}\frac{(2n-2k+j)b^{2n-k}_{k-j}}{2(k-j+1)}y^{2n-2k+j-2}x^{k-j+1}$
-as wished, and this completes the recurrence. In particular, replacing $k=n-1$ in (16), we obtain:
-$\varphi^{n+1}=\frac{1\cdot 3\cdots(2n+1)}{2^{n-1}\cdot n!}ayx^{n}+\sum_{j=0}^{n-1}b^{n+1}_{n-1-j}y^{j+2}x^{n-1-j}\,.$
-We solve the equation (15) for $k=n$ and we get:
-$-2y\varphi^{n}_{x}$ $=\varphi^{n+1}_{y}\,,$
-(17) $-2y\varphi^{n}_{x}$ $=\frac{1\cdot 3\cdots(2n+1)}{2^{n-1}\cdot n!}ax^{n}+\sum_{j=0}^{n-1}(j+2)b^{n+1}_{n-1-j}y^{j+1}x^{n-1-j}\,.$
-From the equality (17), we deduce that $a=0$ and as consequence from this we obtain that $\varphi^{2n+1}=0$. Therefore,
-$\varphi=z\varphi^{2n}+z^{2}\varphi^{2n-1}+\cdots+z^{2n}\varphi^{1}+\varphi^{0}z^{2n+1}=z\psi\,,$
-where $\psi\in F[x,y,z]$ is a homogeneous polynomial of degree $2n$, and as in the previous case, we obtain that $\psi$ is also a solution. ∎
+$\phi^{2n-k-1}_{x}=\frac{P^{2n+1}_{k}(2n-2k-1)}{2}\,ay^{2n-2k-3}x^{k+1}+\sum_{j=0}^{k}\frac{(2n-2k+j)b^{2n-k}_{k-j}}{2}y^{2n-2k+j-2}x^{k-j}\,.$
+Integrating with respect to $x$, we find a constant $b^{2n-k-1}_{0}\in F$ such that $\phi^{2n-k-1}$ is equal to
+$P^{2n+1}_{k+1}ay^{2n-2k-3}x^{k+2}+\sum_{j=0}^{k}\frac{(2n-2k+j)b^{2n-k}_{k-j}}{2(k-j+1)}y^{2n-2k+j-2}x^{k-j+1}+b^{2n-k-1}_{0}y^{2n-k-1}\,$,
+taking into account that $P^{2n+1}_{k+1}=\frac{P^{2n+1}_{k}\cdot(2n-2k-1)}{2\cdot(k+2)}$. Defining
+$b^{2n-k-1}_{k+1-j}=\frac{(2n-2k+j)b^{2n-k}_{k-j}}{2(k-j+1)}$
+for all $0\leq j\leq k$, we get
+$\phi^{2n-k-1}=P^{2n+1}_{k+1}\,ay^{2n-2k-3}x^{k+2}+\sum_{j=0}^{k+1}b^{2n-k-1}_{k+1-j}y^{2n-2k+j-2}x^{k-j+1}$
+as wished, and this completes the recurrence. In particular, replacing k=n1k=n-1 in (18), we obtain:
+$\phi^{n+1}=\frac{1\cdot 3\cdots(2n+1)}{2^{n}\cdot n!}ayx^{n}+\sum_{j=0}^{n-1}b^{n+1}_{n-1-j}y^{j+2}x^{n-1-j}\,.$
+Considering the equation (17) for $k=n$, we get $2y\phi^{n}_{x}=\phi^{n+1}_{y}$, hence
+(19) $2y\phi^{n}_{x}$ $=\frac{1\cdot 3\cdots(2n+1)}{2^{n-1}\cdot n!}ax^{n}+\sum_{j=0}^{n-1}(j+2)b^{n+1}_{n-1-j}y^{j+1}x^{n-1-j}\,.$
+From this equality, we deduce that $a=0$, and consequently $\phi^{2n+1}=0$. Therefore,
+$\varphi=z\phi^{2n}+z^{2}\phi^{2n-1}+\cdots+z^{2n}\phi^{1}+\phi^{0}z^{2n+1}=z\psi\,$,
+where $\psi\in F[x,y,z]$ is a homogeneous polynomial of degree $2n$. Finally, arguing as in item ($a$), we deduce that $\psi$ is also a solution.
 4. Proof of the main results
+We are now ready to prove the main theorems.
 Proof of Theorem 1.
 Suppose that $\varphi=p(xz+y^{2},z)$, where $p$ is a polynomial in $F[t_{1},t_{2}]$. Replacing the partial derivatives
-$\varphi_{x}$ $=z\cdot p_{s}(xz+y^{2},z)\,,$
-$\varphi_{y}$ $=2y\cdot p_{s}(xz+y^{2},z)\,.$
-in the linear PDE (5), one verifies that $\varphi=p(xz+y^{2},z)$ is a solution.
+$\varphi_{x}$ $=z\cdot p_{t_{1}}(xz+y^{2},z)\,$,
+$\varphi_{y}$ $=2y\cdot p_{t_{1}}(xz+y^{2},z)\,.$
+in the linear PDE (7), one readily verifies that $\varphi=p(xz+y^{2},z)$ is a solution.
 Reciprocally, suppose that $\varphi\in F[x,y,z]$ is a solution of the equation (7).
@@
 Indeed, applying Proposition 4 for $n=1$, we obtain:
-$\varphi=\left\{\begin{array}[]{ccl}a(xz+y^{2})+z\psi&\mbox{if}&\deg\varphi=2\,,\\ &&\\ z\psi_{1}&\mbox{if}&\deg\varphi=1\,,\end{array}\right.$
+[…]
 where $\psi$ and $\psi_{1}$ are homogeneous polynomials with $\deg\psi=1$ and $\deg\psi_{1}=0$.
 Then there are constants $b,c\in F$ such that $\psi=cz$ and $\psi_{1}=b\in F$, so (20) is true for $n=1$.
-Let us suppose that (18) holds for homogeneous polynomial solutions of degree $2n$ or $12n-$ 1.
+Let us suppose that (20) holds for homogeneous polynomial solutions of degree $2n$ or 2n12n-1.
 If $\deg\varphi=2n+2$, then by Proposition 4 ($a$), we obtain that $\varphi=a(xz+y^{2})^{2n+2}+z\psi$, where $\psi$ is a solution of the equation (7).
@@
 $=a(xz+y^{2})^{2n+2}+z\left(\sum_{k=1}^{n}a_{2k+1}z^{2k-1}(xz+y^{2})^{2n-2k}\right)\,.$
-If $\deg\varphi=2n+1$, then by Proposition 4 ($b$), we obtain that $\varphi=z\psi$, where $\psi$ is a solution of the equation (5) with $\deg\psi=2n$. By the inductive hypothesis, one has
+Therefore, $\varphi$ has the form $\varphi=p(xz+y^{2},z)$, as wished.
+On the other hand, if $\deg\varphi=2n+1$, then by Proposition 4 ($b$), we obtain that $\varphi=z\psi$, where $\psi$ is a solution of the equation (7) with $\deg\psi=2n$.
+By the inductive hypothesis, one has
 $\psi=\sum_{k=0}^{n}a_{2k}z^{2k}(xz+y^{2})^{2n-2k}\,.$
@@
 $=\sum_{k=1}^{n+1}a_{2k-2}z^{2k-1}(xz+y^{2})^{2(n+1)-2k}\,$,
-which completes the proof. ∎
-Proof of the Theorem 2.
+thus $\varphi$ has the form $\varphi=p(xz+y^{2},z)$, and this completes the proof. ∎
+Proof of the Corollary 3.
 One implication is straightforward.
-Suppose that $\det J(f,g,h)$ is a nonzero constant.
+Suppose that the determinant of $J(f,g,h)$ is a nonzero constant.
 By Proposition 1, $\varphi$ satisfies the linear PDE (7).
-By Theorem 1, the solution of the equation (5) has the form $\varphi=p(xz+y^{2},z)$ where $p$ is a polynomial in $F[t_{1},t_{2}]$.
+By Theorem 1, the solution of the equation (7) has the form $\varphi=p(xz+y^{2},z)$, where $p$ is a polynomial in $F[t_{1},t_{2}]$.
 Finally, by Proposition 2, $(f,g,h)$ is an automorphism and its inverse is given in (3).
@@
 5. Conclusions
-Theorem 2 says that the automorphisms of the form (2) satisfy the well know Jacobian conjecture via the approach of solving first order partial differential equations.
-A further problem is whether those automorphisms might be naturally extended on polynomial rings with more than three variables satisfying the aforementioned result.
+In Corollary 3, we show that the Nagata automorphisms of the form (2) satisfy the well-known Jacobian conjecture.
+The proof is based on algebraic techniques for solving first-order linear partial differential equations.
+This result suggests a new strategy for approaching this longstanding open problem.